OpenStudy (anonymous):

A paint can has a diameter of 17 centimeters and a height of 24 centimeters. How many cubic centimeters of paint will fill the can?

OpenStudy (ghazi):

v= pi*r^2*l=3.14*(8.5)^2*24

OpenStudy (ghazi):

\[3.14*(8.5)^2*24\]